# Bridging the performance-gap between target-free and target-based reinforcement learning

### v2

@@ 5.4 Scaling Up to Language Models
In this experiment, we evaluate iS-QN on an offline RL language processing task. Specifically, we focus on Implicit Language $Q$-Learning (ILQL, Snell et al. (2023)), a method introduced with a target network. It adapts implicit $Q$-learning (Kostrikov et al., 2023) to the language domain by sampling action tokens from a policy, learned with supervised learning, and weighted by the advantage computed from the $Q$-function. We evaluate ILQL on the Wordle game (Lokshtanov & Subercaseaux, 2022), a multi-turn game where the agent guesses a hidden word and receives feedback after each attempt. As in Snell et al. (2023), we choose the GPT-$2$ small architecture, which results in TB-ILQL using $264$ million parameters. In Figure 7 (left), we note that while a performance drop is noticeable, the target-free approach does not perform significantly worse than the target-based variant. Importantly, sharing parameters and learning $K=9$ Bellman iterations in parallel improves the learning speed of the target-free approach by $10\%$ without significantly increasing the memory footprint. This leads iS-QN to save $88$ million parameters compared to the original approach.
@@ 6 Limitation and Conclusion
The proposed approach introduces the number of Bellman updates $K$ to learn in parallel as a new hyperparameter and there seems to be a different optimal value for each setting. However, we observe that none of the values lead to a dramatic performance drop. In this work, we focus on reducing the memory footprint of the function approximators. Depending on the setting, other objects such as the replay buffer and the optimizer can occupy a large portion of the RAM. We remark that the proposed approach can be combined with other works addressing these issues (Vasan et al., 2024). Additionally, the proposed approach reduces the memory footprint during training but uses the same amount during inference, which is complementary to pruning methods that use more memory during training and less during inference (Graesser et al., 2022). As reported in Figure 10, iS-QN does not reduce the training time or the number of floating-point operations, except for the language processing task for which the temporal-difference error can be computed with a single pass through the network.
We introduced a simple yet efficient method for mitigating the performance drop that occurs when removing the target network in deep value-based reinforcement learning, while maintaining a low memory footprint. This is made possible by storing a copy of the last linear layer of the online network and using the features of the online network as input to this frozen linear head to construct the regression target. From there, more heads can be added to learn multiple Bellman iterations in parallel. We demonstrated that this new algorithm, iterated Shared $Q$-Networks, improves over the target-free approach and yields higher returns when the number of heads increases. We believe that combining iS-QN with pruning and/or quantization methods is a promising direction for future work to facilitate online learning on resource-constrained settings, without sacrificing performance.
@@
Special care was taken to ensure this work is reproducible. The code will be made open source upon acceptance and is shared in the supplementary material. It contains the list of dependencies and their exact version that was used to generate the results. To ease reproducibility, all hyperparameters are listed in Appendix D, and the individual training curves are shown in Appendix E.
@@ Appendix A Experiment Setup
Atari games selection  Our evaluations on the CNN architecture were performed on the $15$ games recommended by Graesser et al. (2022). They were chosen for their diversity of Human-normalized score that DQN reaches after being trained on $200$ million frames, as shown in Figure 9. As the IMPALA architecture increases the training length, we removed $5$ games, while maintaining diversity in the final scores to reduce the computational budget. For the offline experiment, we used the datasets provided by Gulcehre et al. (2020). As the game Tutankham is not available in the released dataset, we replaced it with Qbert, indicated with an asterisk in Figure 9.
@@
Wordle setup  Our codebase is a fork of the repository shared by the authors (Snell et al., 2023), from which we implemented the target-free and the iterated Shared Features approaches. We refer to the original paper for extensive details about the setup.
@@ Appendix C Algorithmic Details
Aggregating individual losses  In Equation 1, we define the loss of iS-QN as the sum of losses over each Bellman iteration. Other ways of aggregating the losses are possible. Nonetheless, we decided to stick to the version proposed by Vincent et al. (2025) and leave this investigation for future work. We provide a first alternative in Section 5.3 that provides a performance boost by discounting the following terms by a factor of $0.25$. While it is true that taking the sum of temporal differences increases the magnitude of the loss, it has a different impact on the updates than simply multiplying the learning rate by the number of terms in the loss. Indeed, the Adam optimizer (Kingma & Ba, 2015) first normalizes the gradient with a running statistic before applying the learning rate. Therefore, changing the aggregation mechanism has a greater impact on the direction of the update than on its magnitude. This is why we do not compare iS-QN against baselines instantiated with different learning rates.
Sampling actions  Following Vincent et al. (2025), at each environment interaction, an action is sampled from a single head chosen uniformly as shown in Line 3 in Algorithm 1. The authors motivate this choice by arguing that it allows each $Q$-function to interact with the environment, thereby avoiding passive learning, identified by Ostrovski et al. (2021). This choice is further justified by an ablation study (see Figure $19$ in Vincent et al. (2025)) demonstrating a stronger performance against another sampling strategy consisting of sampling one head for each episode, as proposed in Osband et al. (2016).
@@
Appendix D List of Hyperparameters

### v3

@@ 5.4 Scaling Up to Language Models
In this experiment, we evaluate iS-QN on an offline RL language processing task. Specifically, we focus on Implicit Language $Q$-Learning (ILQL, Snell et al. (2023)), a method introduced with a target network. It adapts implicit $Q$-learning (Kostrikov et al., 2023) to the language domain by sampling action tokens from a policy, learned with supervised learning, and weighted by the advantage computed from the $Q$-function. We evaluate ILQL on the Wordle game (Lokshtanov and Subercaseaux, 2022), a multi-turn game where the agent guesses a hidden word and receives feedback after each attempt. As in Snell et al. (2023), we choose the GPT-$2$ small architecture, which results in TB-ILQL using $264$ million parameters. In Figure 7 (left), we note that while a performance drop is noticeable, the target-free approach does not perform significantly worse than the target-based variant. Importantly, sharing parameters and learning $K=9$ Bellman iterations in parallel improves the learning speed of the target-free approach by $10\%$ without significantly increasing the memory footprint. This leads iS-QN to save $88$ million parameters compared to the original approach.
Figure 8: Increasing the learning speed in a streaming scenario on $7$ Atari games with the CNN architecture and LayerNorm (LN). iS-Stream Q($\lambda$) $K=3$ improves over the target-free approach by more than $10\%$, outperforming or performing on par with the baseline on 66 out of the $7$ games.
5.5 Streaming Reinforcement Learning
We consider the streaming setting in which the RL agent only learns from a stream of data. In this setting, the agent does not have access to a replay buffer, batch updates, or parallel environments. These constraints position the RL agent in a drastically different setting from the previously considered ones, as replay buffers and batch updates greatly stabilize the learning process (Vasan et al., 2024). Elsayed et al. (2024) introduced Stream Q($\lambda$), an algorithm adapted for the streaming setting. It is an adaptation of the original Q($\lambda$) algorithm that uses observation and reward normalization, sparse initialization, layer norm, and an adaptive step-size optimizer. In this experiment, we combine Stream Q($\lambda$) with the presented approach to obtain iS-Stream Q($\lambda$). We remark that Stream Q($\lambda$) was introduced without a target network; therefore, we focus only on this version (TF-Stream Q($\lambda$)). We fix a target update period of $T=10$ for iS-Stream Q($\lambda$). In Figure 8, we report the performance of the target-free approach and the presented approach for three values of $K$ across $7$ Atari games that were selected for their diversity in human-normalized score (see Figure 10). iS-Stream Q($\lambda$) $K=1$ performs similarly to the target-free variant. Remarkably, increasing the number of heads to 44 ($K=3$) improves the learning speed compared to the target-free approach. In this setting, increasing the number of heads until $9$ Bellman iterations are learned in parallel is not helpful. The absence of batch updates increases the variance of the gradient steps, potentially leading to unstable updates for the shared parameters for a large number of heads.
@@ 6 Limitation and Conclusion
The proposed approach introduces the number of Bellman updates $K$ to learn in parallel as a new hyperparameter and there seems to be a different optimal value for each setting. However, we usually observe a stable increase in performance as $K$ grows until the benefit of the approach starts to diminish. Therefore, we recommend increasing $K$ until performance begins to decrease, as is commonly done with the learning rate. In Sections D.2, D.3, and D.4, we provide an extensive analysis of the dependency of iS-QN on its hyperparameters. In this work, we focus on reducing the memory footprint of the function approximators. Depending on the setting, other objects such as the replay buffer and the optimizer can occupy a large portion of the RAM. We remark that the proposed approach can be combined with other works addressing these issues (Vasan et al., 2024). Additionally, the proposed approach reduces the memory footprint during training but uses the same amount during inference, which is complementary to pruning methods that use more memory during training and less during inference (Graesser et al., 2022). As reported in Figure 11, iS-QN does not reduce the training time or the number of floating-point operations, except for the language processing task for which the temporal-difference error can be computed with a single pass through the network.
We introduced a simple yet efficient method for mitigating the performance drop that occurs when removing the target network in deep value-based reinforcement learning, while maintaining a low memory footprint. This is made possible by storing a copy of the last linear layer of the online network and using the features of the online network as input to this frozen linear head to construct the regression target. From there, more heads can be added to learn multiple Bellman iterations in parallel. We demonstrated that this new algorithm, iterated Shared $Q$-Networks, improves over the target-free approach and yields higher returns when the number of heads increases. We believe that combining iS-QN with mixed-precision training methods is a promising direction for future work to facilitate online learning in resource-constrained settings without sacrificing performance. In Section D.5, we provide a pilot study demonstrating positive results.
Acknowledgments
This research was supported by “Third Wave of AI”, funded by the Excellence Program of the Hessian Ministry of Higher Education, Science, Research and Art, and by the grant “Einrichtung eines Labors des Deutschen Forschungszentrum für Künstliche Intelligenz (DFKI) an der Technischen Universität Darmstadt”. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b187cb. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. This work was also partially supported by the German Federal Ministry of Research, Technology and Space (BMFTR) under the Robotics Institute Germany (RIG).
@@ Acknowledgments
Special care was taken to ensure this work is reproducible. The code is available at https://github.com/theovincent/iS-DQN and is shared in the supplementary material. It contains the list of dependencies and their exact version that was used to generate the results. To ease reproducibility, all hyperparameters are listed in Appendix E, and the individual training curves are shown in Appendix F.
@@
Carbon Impact
As recommended by Lannelongue and Inouye (2023), we used GreenAlgorithms (Lannelongue et al., 2021) and ML $CO_{2}$ Impact (Lacoste et al., 2019) to compute the carbon emission related to the production of the electricity used for the computations of our experiments. We only consider the energy used to generate the figures presented in this work and ignore the energy used for preliminary studies and for building the computing infrastructure. The estimations vary between $2.63$ and $2.88$ tonnes of $\text{CO}_{2}$ equivalent. As a reminder, the Intergovernmental Panel on Climate Change advocates a carbon budget of $2$ tonnes of $\text{CO}_{2}$ equivalent per year per person.
@@ Appendix A Experiment Setup
Atari games selection  Our evaluations on the CNN architecture were performed on the $15$ games recommended by Graesser et al. (2022). They were chosen for their diversity of Human-normalized score that DQN reaches after being trained on $200$ million frames, as shown in Figure 10. As the IMPALA architecture increases the training length, we removed $5$ games, while maintaining diversity in the final scores to reduce the computational budget. For the offline experiment, we used the datasets provided by Gulcehre et al. (2020). As the game Tutankham is not available in the released dataset, we replaced it with Qbert, indicated with an asterisk in Figure 10. The codebase is adapted from the code released by Vincent et al. (2025a).
@@
Wordle setup  Our codebase is a fork of the repository shared by the authors (Snell et al., 2023), from which we implemented the target-free and the iterated Shared Features approaches. We refer to the original paper for extensive details about the setup. Every algorithm was given a budget of 600000600~000 gradient steps. This value was chosen by running TB-ILQL until the performance reported in Table $6$ of Snell et al. (2023) was reached.
Streaming RL setup  We follow the implementation choices made by Elsayed et al. (2024). We detail here the differences with the classical Atari setup described earlier. Namely, the source of stochasticity is replaced from sticky actions to a random number of NO OP actions chosen uniformly between 0 and $30$. The life loss signal is used for episode termination instead of the game over signal. The last hidden linear layer is composed of $256$ neurons instead of $512$. The neural network weights are initialized with a sparsity level of $90\%$. The optimizer’s hyperparameters are $\kappa=2$ and the default learning rate is $1$. The trace coefficient is set to $\lambda=0.8$.
@@ Appendix B Training Time and Floating-point Operations
We point out that when computing the algorithm’s runtime, the environment steps are simulated, which does not reflect the real-world duration. For example, an episode of an Atari game can be executed significantly faster in simulation than in real life. This means that if real-world durations were taken into account, sample-efficient algorithms would achieve better performance earlier than less sample-efficient methods.
@@ Appendix C Algorithmic Details
Aggregating individual losses  In Equation 1, we define the loss of iS-QN as the sum of losses over each Bellman iteration. Other ways of aggregating the losses are possible. Nonetheless, we decided to stick to the version proposed by Vincent et al. (2025b) and leave this investigation for future work. We provide a first alternative in Section 5.3 that provides a performance boost by discounting the following terms by a factor of $0.25$. While it is true that taking the sum of temporal differences increases the magnitude of the loss, it has a different impact on the updates than simply multiplying the learning rate by the number of terms in the loss. Indeed, the Adam optimizer (Kingma and Ba, 2015) first normalizes the gradient with a running statistic before applying the learning rate. Therefore, changing the aggregation mechanism has a greater impact on the direction of the update than on its magnitude. This is why we do not compare iS-QN against baselines instantiated with different learning rates.
Sampling actions  Following Vincent et al. (2025b), at each environment interaction, an action is sampled from a single head chosen uniformly as shown in Line 3 in Algorithm 1. The authors motivate this choice by arguing that it allows each $Q$-function to interact with the environment, thereby avoiding passive learning, identified by Ostrovski et al. (2021). This choice is further justified by an ablation study (see Figure $19$ in Vincent et al. (2025b)) demonstrating a stronger performance against another sampling strategy consisting of sampling one head for each episode, as proposed in Osband et al. (2016).
@@
Appendix D Additional Experiments
In this section, we train each agent for 4040M frames rather than 100100M to reduce the computational budget.
D.1 Comparison with MellowMax DQN
Figure 12: When using the MellowMax operator, the presented approach (iS-Mellow) effectively reduces the performance gap between the target-free approach (TF-Mellow) and the target-based approach (TB-Mellow).
We combine the MellowMax operator (Kim et al., 2019) with the presented idea. We select the games that were presented in the original paper (Kim et al., 2019), and use the same temperature coefficients ($\omega=1\ 000$ for Breakout, and $\omega=30$ for Seaquest). Remarkably, in Figure 12, the performance gap between the target-free and target-based approaches is reduced by the presented approach.
D.2 Ablation Study on the Target Update Period
Figure 13: Ablation study on the target update period $T$ for the game Tutankham. iS-DQN’s performance remains stable when varying the target update period. This shows that it is not sensitive to the target update period for this game.
We perform an ablation study on the frequency at which the heads of the iS-DQN agent are updated. This hyperparameter plays a role similar to the target update period $T$ in DQN, allowing both algorithms to propagate rewards, as explained in Section 4. In Figure 13, we report the performance of the target-based approach and the presented approach for $5$ different values of $T$ on the game Tutankham. As a reminder, $T=8000$ is the default value used in Section 5. We also report the performance of the target-free variant as a horizontal line since this algorithm does not use this hyperparameter. Interestingly, the performance of iS-DQN remains similar across different values of $T$, demonstrating its robustness to this hyperparameter in this game. Remarkably, iS-DQN $K=9$ outperforms both the target-free and target-based variants across all target update periods.
D.3 Ablation Study on the Number of Bellman Updates
Figure 14: Ablation study on the number of Bellman updates learned in parallel $K$ and the size of the shared features. iS-DQN achieves similar performance for large values of $K$ and consistently outperforms the target-free approach.
We analyze the sensitivity of iS-DQN with respect to the number of Bellman updates $K$ learned in parallel. As discussed in Section 5.1, the richness of the shared features can affect the behavior of the method. To account for this aspect, we vary the size of the last linear layer from half the default value ($256$ neurons, "Width $\times~0.5$") to twice the default ($1024$ neurons, "Width $\times~2.0$"). To perform this study, we choose the games CrazyClimber and Tutankham as the performance of iS-DQN on those games for the default setting ($512$ neurons, "Width $\times~1.0$") reflects the overall performance presented in Figure 3 on the $15$ games. In Figure 14, we report the performance of iS-DQN for the three different settings with the number of heads varying from 22 ($K=1$) to $50$ ($K=49$). As expected, the performance of iS-DQN is similar for large values of $K$ ($K\geq 9$). Importantly, every value of $K$ results in improved performance compared to the target-free variant, even if a slight decrease in performance occurs at the largest values of $K$. This is expected, as the potential of learning multiple Bellman updates in parallel can be better realized with the higher representational capacity.
D.4 Automatic Tuning of the Importance of Each Bellman Update
In this section, we explore a way of tuning the importance of each Bellman update during training. For that, we propose weighting each term in the loss with a different coefficient. We note the coefficient $(\alpha_{k})_{k=1}^{K}$. Intuitively, more weight should be given to a Bellman update that produces a gradient aligned with the gradients of the other Bellman updates. We will show that meta-learning those coefficients using the concept of meta-gradient reinforcement learning (Xu et al., 2018) provides a natural way to achieve this goal. Indeed, the update rule coming from the meta-gradient algorithm of a coefficient $\alpha_{k}$ linked to the $k^{\text{th}}$ Bellman update, depends on the dot product between the gradient of the loss of the $k^{\text{th}}$ Bellman update and the gradient of the sum of all Bellman updates:
[…]
where $d=(s,a,r,s^{\prime})$ is a sample, $\alpha$ is the vector of coefficients $(\alpha_{k})_{k=1}^{K}$, $\theta(\alpha)$ is the parameters after one gradient step starting from $\theta$, $\omega$ is the shared parameters, $\omega_{k}$ is the parameters of the $k^{\text{th}}$ head such that $\theta_{k}=(\omega,\omega_{k})$, $\lambda_{\theta}$ is the learning rate of the parameters, and $\lambda_{\alpha}$ is the meta learning rate.
In the following, we define the meta-optimization problem, analyse the empirical results, and finally derive Equation D.4. To define the meta-optimization problem, we first stress the dependence of the learned parameters on the meta-parameters by noting them as a function of $\alpha$: $\theta^{*}(\alpha)$. Therefore, the inner loop of the meta-gradient optimization problem is defined as $\theta^{*}(\alpha)=\min_{\theta}\sum_{k=1}^{K}\alpha_{k}\mathcal{L}^{\text{QN}}_{d}(\theta_{k},\theta_{k-1})$. This leads to the following optimization problem:
$\min_{\alpha}\sum_{k=1}^{K}\mathcal{L}^{\text{QN}}_{d}(\theta^{*}(\alpha)_{k},\theta^{*}(\alpha)_{k-1})\quad\text{s.t. }\theta^{*}(\alpha)=\min_{\theta}\sum_{k=1}^{K}\alpha_{k}\mathcal{L}^{\text{QN}}_{d}(\theta_{k},\theta_{k-1})$
Figure 15: Top: Evolution of the meta-learned coefficients for each Bellman update during training for $3$ Atari games. Remarkably, the coefficients converge to similar values, indicating that setting equal weights for each Bellman update is a good static choice for this setting. Bottom: Meta iS-DQN $K=9$ performs on par with iS-DQN $K=9$, which is coherent with the values of the learned coefficients.
We evaluate this approach for $K=9$ on $3$ Atari games, selected for their diversity in human-normalized scores (see Figure 10). We parameterize the meta-parameters $(\alpha_{k})_{k=1}^{K}$ as logits $(z_{k})_{k=1}^{K}$ and apply the softmax function such that the loss is a convex combination of the individual terms. In Figure 15 (top), we report the learned meta-parameters, which we use as a proxy for the importance of each Bellman update during training. Remarkably, after 1010M frames, the coefficients converge to identical values, giving equal importance to each Bellman update. This indicates that setting equal weights for each Bellman update is a good static choice for this setting, as proposed in Equation 1. This explains why in Figure 15 (bottom), Meta iS-DQN $K=9$ performs on par with iS-DQN $K=9$, as iS-DQN $K=9$ uses equal static weights during training. Importantly, Meta iS-DQN $K=9$ performs better than iS-DQN $K=1$.
To derive Equation D.4, we assume that the meta-parameters only influence one update step, as in Xu et al. (2018). We also simplify the computations by choosing SGD as the optimizer instead of Adam, noting $\theta(\alpha)$ the parameters after one gradient step, starting from $\theta$. The meta-parameter update, for a sample $d=(s,a,r,s^{\prime})$ is
$\alpha_{k}\leftarrow\alpha_{k}-\lambda_{\alpha}\nabla_{\alpha_{k}}\sum_{i=1}^{K}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},$ $\theta(\alpha)_{i-1})=\alpha_{k}-\lambda_{\alpha}\sum_{i=1}^{K}\nabla_{\alpha_{k}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1}),$ (3)
$\text{where }\nabla_{\alpha_{k}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})$ $=\nabla_{\theta(\alpha)_{i}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\nabla_{\alpha_{k}}\theta(\alpha)_{i}$
$+\nabla_{\theta(\alpha)_{i-1}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\nabla_{\alpha_{k}}\theta(\alpha)_{i-1}$ (4)
Equation 4 comes from the chain rule. We remark that for any $k$, $\nabla_{\theta_{k-1}}\mathcal{L}_{d}^{\text{QN}}(\theta_{k},\theta_{k-1})=0$, since the target network is frozen. We split the parameters of the heads from the shared parameters as in Section 4: $\theta(\alpha)_{i}=(\omega(\alpha),\omega(\alpha)_{i})$, where $\omega$ represents the shared parameters and $\omega_{i}$ the parameters of the $i^{\text{th}}$ head. We now focus on the gradient of the $i^{\text{th}}$ head parameters with respect to $\alpha_{k}$:
$\nabla_{\alpha_{k}}\omega(\alpha)_{i}$ $\stackrel{{\scriptstyle(a)}}{{=}}\nabla_{\alpha_{k}}\left(\omega_{i}-\lambda_{\theta}\nabla_{\omega_{i}}\sum_{j=1}^{K}\alpha_{j}\mathcal{L}_{d}^{\text{QN}}(\theta_{j},\theta_{j-1})\right)\stackrel{{\scriptstyle(b)}}{{=}}-\lambda_{\theta}\nabla_{\alpha_{k}}\left(\sum_{j=1}^{K}\alpha_{j}\nabla_{\omega_{i}}\mathcal{L}_{d}^{\text{QN}}(\theta_{j},\theta_{j-1})\right)$
$\stackrel{{\scriptstyle(c)}}{{=}}-\lambda_{\theta}\nabla_{\alpha_{k}}\left(\alpha_{i}\nabla_{\omega_{i}}\mathcal{L}_{d}^{\text{QN}}(\theta_{i},\theta_{i-1})\right)\stackrel{{\scriptstyle(d)}}{{=}}-\lambda_{\theta}\mathbbm{1}_{k=i}\nabla_{\omega_{i}}\mathcal{L}_{d}^{\text{QN}}(\theta_{i},\theta_{i-1}).$ (5)
$(a)$ comes from the definition of $\omega(\alpha)_{i}$, $(b)$ uses the fact that $\omega_{i}$ and $\lambda_{\theta}$ do not depend on $\alpha_{k}$, $(c)$ uses the fact that the parameters of the $i^{\text{th}}$ head do not influence the parameters of the other heads, and $(d)$ uses the fact that $\nabla_{\alpha_{k}}\alpha_{i}=\mathbbm{1}_{k=i}$. Similarly, we derive the gradient of the shared parameters $\omega(\alpha)$ with respect to $\alpha_{k}$:
$\nabla_{\alpha_{k}}\omega(\alpha)$ $\stackrel{{\scriptstyle(e)}}{{=}}\nabla_{\alpha_{k}}\left(\omega-\lambda_{\theta}\nabla_{\omega}\sum_{j=1}^{K}\alpha_{j}\mathcal{L}_{d}^{\text{QN}}(\theta_{j},\theta_{j-1})\right)\stackrel{{\scriptstyle(f)}}{{=}}-\lambda_{\theta}\nabla_{\alpha_{k}}\left(\sum_{j=1}^{K}\alpha_{j}\nabla_{\omega}\mathcal{L}_{d}^{\text{QN}}(\theta_{j},\theta_{j-1})\right)$
$\stackrel{{\scriptstyle(g)}}{{=}}-\lambda_{\theta}\left(\sum_{j=1}^{K}\nabla_{\alpha_{k}}\alpha_{j}\nabla_{\omega}\mathcal{L}_{d}^{\text{QN}}(\theta_{j},\theta_{j-1})\right)\stackrel{{\scriptstyle(h)}}{{=}}-\lambda_{\theta}\left(\sum_{j=1}^{K}\mathbbm{1}_{k=j}\nabla_{\omega}\mathcal{L}_{d}^{\text{QN}}(\theta_{j},\theta_{j-1})\right)$
$\stackrel{{\scriptstyle(i)}}{{=}}-\lambda_{\theta}\nabla_{\omega}\mathcal{L}_{d}^{\text{QN}}(\theta_{k},\theta_{k-1}).$ (6)
$(e)$ comes from the definition of $\omega(\alpha)$, $(f)$ uses the fact that $\omega$ and $\lambda_{\theta}$ do not depend on $\alpha_{k}$, and the linearity of the sum, $(g)$ uses the linearity of the gradient operator, $(h)$ uses the fact that $\nabla_{\alpha_{k}}\alpha_{j}=\mathbbm{1}_{k=j}$, and in $(i)$, we evaluate the indicator function. Plugging Equations D.4 and D.4 in Equation 4, we obtain:
$\nabla_{\alpha_{k}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})\stackrel{{\scriptstyle(j)}}{{=}}\nabla_{\theta(\alpha)_{i}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\nabla_{\alpha_{k}}\theta(\alpha)_{i}$
$\stackrel{{\scriptstyle(k)}}{{=}}\nabla_{\omega(\alpha)_{i}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\nabla_{\alpha_{k}}\omega(\alpha)_{i}+\nabla_{\omega(\alpha)}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\nabla_{\alpha_{k}}\omega(\alpha)$
$\stackrel{{\scriptstyle(l)}}{{=}}-\lambda_{\theta}\nabla_{\omega(\alpha)_{i}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\mathbbm{1}_{k=i}\nabla_{\omega_{i}}\mathcal{L}_{d}^{\text{QN}}(\theta_{i},\theta_{i-1})$
$\quad-\lambda_{\theta}\nabla_{\omega(\alpha)}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\nabla_{\omega}\mathcal{L}_{d}^{\text{QN}}(\theta_{k},\theta_{k-1})$ (7)
$(j)$ comes from Equation 4, $(k)$ uses the decomposition between the head parameters and the shared parameters, and $(l)$ comes from Equations D.4 and D.4. Finally, plugging Equation D.4 in Equation 3 and using the linearity of the dot product, we obtain the result presented in Equation D.4:
$\alpha_{k}\leftarrow$ $\alpha_{k}+\lambda_{\alpha}\lambda_{\theta}\nabla_{\omega(\alpha)_{k}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{k},\theta(\alpha)_{k-1})^{T}\nabla_{\omega_{k}}\mathcal{L}_{d}^{\text{QN}}(\theta_{k},\theta_{k-1})$
$\quad+\lambda_{\alpha}\lambda_{\theta}\sum_{i=1}^{K}\nabla_{\omega(\alpha)}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})^{T}\nabla_{\omega}\mathcal{L}_{d}^{\text{QN}}(\theta_{k},\theta_{k-1}).$
$=\alpha_{k}+\lambda_{\alpha}\lambda_{\theta}\nabla_{\omega(\alpha)_{k}}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{k},\theta(\alpha)_{k-1})^{T}\nabla_{\omega_{k}}\mathcal{L}_{d}^{\text{QN}}(\theta_{k},\theta_{k-1})$
$\quad+\lambda_{\alpha}\lambda_{\theta}\nabla_{\omega(\alpha)}\left[\sum_{i=1}^{K}\mathcal{L}^{\text{QN}}_{d}(\theta(\alpha)_{i},\theta(\alpha)_{i-1})\right]^{T}\nabla_{\omega}\mathcal{L}_{d}^{\text{QN}}(\theta_{k},\theta_{k-1}).$
D.5 Pilot Study: Mixed Precision Training of iS-QN
Figure 16: When reducing the precision of the operations performed during the forward and backward passes of the neural network from float3232 to float1616, iS-QN still bridges the performance gap between the target-free and target-based approaches.
To further reduce the resources required during training, the presented approach can be combined with other techniques that also aim at reducing resource requirements. We provide an initial study that combines the idea of performing operations with lower precision to further reduce the memory footprint (Micikevicius et al., 2018) with the presented approach. Specifically, the computations during the forward and backward passes of the neural network are performed in float1616b rather than float3232. In Figure 16, we show that the performance gap between the target-free and target-based approaches remains when the operations are performed with lower precision. Remarkably, iS-DQN $K=1$ closes the performance gap, and iS-DQN $K=9$ further boosts the learning speed of the target-free approach, outperforming the target-based approach when the precision of the operations is reduced. Overall, we observe a slight decrease in performance across all algorithms, which is expected, as the precision of the operations during the forward and backward passes decreases.
Appendix E List of Hyperparameters